300 JELLY BEANS COST X AMOUNT OF DOLLARS. HOW MANY JELLY BEANS COULD YOU BUY FOR 50 CENTS AT THE SAME RATE?
Answer by stanbon(48551)

300 JELLY BEANS COST X AMOUNT OF DOLLARS. HOW MANY JELLY BEANS COULD YOU BUY FOR 50 CENTS AT THE SAME RATE?
-------------
X dollars = 100X cents
Set up a proportion as follows:
300/(100x cents) = ?/(50 cents)
Multiply both sides by 50 cents to get:
? = (50 cents)[300/100x cents]
Reduce (50 cents)/(100X cents) to get:
? = (1/(2X))300
? = 150/X
You can get 150/X jelly beans for 50 cents.
Cheers,
Stan H.